OpenStudy (anonymous):

integral 2/square root of x dx from 9 to 1

OpenStudy (anonymous):

$$\int_9^1\frac2{\sqrt{x}}dx=2\int_9^1x^{-\frac12}dx=2\left[2x^\frac12\right]_9^1=4\sqrt1-4\sqrt9=4-12=-8$$

OpenStudy (anonymous):

thanks :)